Hi i'm trying to solve for X, a/bx+c/d=e.
I have reduced if to a/bx=ed-c/d but can't see how to get to the answer.
which i belive is x=da/b(ed-c).
Please help
Thank Darryl

Answer by venugopalramana(3286)

Hi i'm trying to solve for X, a/bx+c/d=e.
USE BRACKETS TO SHOW THE THINGS PROPERLY AND TO AVOID MISTAKES.
(a/b)x+(c/d)=e.
I have reduced if to a/bx=ed-c/d
NO...IT IS
(a/b)x=e-(c/d) = (ED-C)/D
SO NOW WE GET
X={(ED-C)}/(A/B)
=B(ED - C )/A......IS THE ANSWER.HOWEVER IF YOUR PROBLEM IS REALLY
(a/bx)+(c/d)=e....THEN
(A/BX)= E - C/D = (ED-C)/D
A/B = X*(ED - C)/D
X = (A/B)/{(ED-C)/D}
X = AD/B(ED - C)WHICH IS THE ANSWER YOU ARE SHOWING BELOW...SO NOW YOU UNDERSTAND THE IMPORTANCE OF PUTTING PROPER BRACKETS.
but can't see how to get to the answer.
which i belive is x=da/b(ed-c).